# The Basis of Foot-Sortable Sock Orderings

### v1

@@ 6. Proofs of Lemmas 26, 27, and 28
Case 1: A sandwich of the form $ded$ blocks $a$ in $X$. Note that since $d\notin R^{\prime}$, we have $d\in S^{\prime}$. So, if $S^{\prime}$ does not contain $dat$ or $adt$, then $S^{\prime}$ must contain $atde$ and $e\neq t$. So to show that $X$ satisfies (i), we only need to show that $e\notin P$. First note that $e\neq a$, since $ded$ blocks $a$. If $e=r$, then, since we always have $art\subseteq P$, the sequence $S^{\prime}dP\subseteq X$ contains $atdrdart\cong abdtdatb\cong C_{2,5}$, a contradiction. We get a similar contradiction if $P=arbrt$ or $P=arcbct$ and $e=b$, since in either case $P$ contains $abt$, so $S^{\prime}dP$ contains $atdbdabt\cong abdtdatb\cong C_{2,5}$. If $P=arcbct$ and $e=c$, then $P$ contains $act$, so $S^{\prime}dP$ contains $atdcdact\cong abdtdatb\cong C_{2,5}$. This completes the proof of this case.
@@
Case 2.1: $e=t$. In this case, $ta=ea\subseteq R^{\prime}$ and, by Claim 37 a sandwich of the form $trt$ does not block $a$ in $X=S^{\prime}dR^{\prime}$. The fact that $at\subseteq S^{\prime}$ then implies that $rta\nsubseteq R^{\prime}$. Therefore, there exists a $t$ in $R^{\prime}$ that is before both an $a$ in $R^{\prime}$ and the first $r$ in $R^{\prime}$. So, if $P=arat$, then we have either $tarat=eP$ or $atrat$ in $R^{\prime}$, so (ii) holds. Furthermore, if $P=rart$, we have $trart=eP$ in $R^{\prime}$, so (ii) holds. If $P=arbrt$, then one option is $tarbrt=eP\subseteq R^{\prime}$, and (ii) holds in this case. By Claim 37, there is no $rbr$ sandwich that blocks $a$, so the only other possibilities when $P=arbrt$ are for $R^{\prime}$ to contain $atrabrt$ or $atrbart$, and these both contain $trart$, so (ii) holds. Finally, assume $P=arcbct$. If $tarcbct=eP\subseteq R^{\prime}$, then $R^{\prime}$ contains $tacbct$, so (ii) holds. Otherwise, because there is no $cbc$ sandwich that blocks $a$, the pattern $R^{\prime}$ must contain one of $atracbct$, $atrcabct$, or $atrcbact$. These contain $tacbct$, $tcact$, and $tcact$, respectively, so (ii) holds.
Case 2.2: $e=r$. In this case, $r\in S^{\prime}$ and $ra=ea\subseteq R^{\prime}$. Note that, by Claim 37, there is no sandwich of the form $rtr$ that blocks $a$ in $X$. Therefore, $r$ must follow $at$ in $S^{\prime}$, so $atr\subseteq S^{\prime}$. Since $ar\subseteq P$ and $t\in P$, either $art$, $atr$, or $tar=tI(r)ar$ is in $P\subseteq R$. Therefore, Lemma 27 implies that a subpattern of $X$ is in $\Gamma$, a contradiction.
@@
Assume first that $e\in R$ and recall that $r$ was defined to be the first color in $R$ that is neither $t$ nor $a$. Therefore, $rea\subseteq R^{\prime}$, which implies that a sandwich of the form $ere$ blocks $a$ in $X\supseteq S^{\prime}dR^{\prime}$. Note that this, with Claim 37, implies that $e\notin P$. Recall that Claim 37 implies that every color in $atP$ is unsortable except $a$. If $e$ is blocked by a sandwich consisting of colors only in $P$, then the pattern formed by removing all colors except those in $P$ and $e$ from $X-d$ is trivially unsortable. This contradicts the assumption that all strict subpatterns of $X$ are sortable. Therefore, we can assume that $e$ is not blocked by a sandwich consisting of colors only in $P$ in $X$.
Now assume that $e\notin R$. In this case, $P\subseteq R$ implies that $e\notin P$. By the definition of $R^{\prime}$, the fact that $e\in R^{\prime}$ implies that the last $e$ follows the last $d$ in $X$. Since $e\notin R$ and the state after $d$ is sorted is $(S,R)$, this means that at some sorting step before the color $d$ was sorted, $de$ appeared on the stack. Therefore, $e$ was sorted before $d$. Since $d$ was sorted before any color in $P$, the color $e$ is not blocked by a sandwich consisting only of colors in $P$ in $X$.
So, in all cases, $e\notin P$ and $e$ is not blocked by a sandwich consisting only of colors in $P$ in $X$. In particular, the pattern $ata$ does not block $e$ in $X$. This means that $e$ does not follow $a$ in $R^{\prime}$. Note first that if $eP\subseteq R^{\prime}$, then (iii) holds. So we can assume $eP\nsubseteq R^{\prime}$. This, with the fact that $e$ does does not follow $a$ in $R^{\prime}$ and $e\in R^{\prime}$, implies that $P=rart$ and $reart\subseteq R^{\prime}$. Therefore, $X=S^{\prime}dR^{\prime}$ contains either $eatdreart$, $aetdreart$, or $atedreart$. These contain $eatrear\cong dabtdat\cong I_{1,0}$, $aetreart\cong adbtdatb\cong\mathcal{C}_{2,4}$, and $atereart\cong abdtdatb\cong\mathcal{C}_{2,5}$, respectively, which is a contradiction. ∎
When Claim 38(i) holds, then $X$ must contain one of the patterns in the following table.
@@
If Claim 38(iii) holds, then by swapping $d$ and $e$, we can see that $X$ must contain one of these same patterns. Finally, if Claim 38(ii) holds, then $X$ must contain one of the patterns in the following table.

### v2

@@ 2.6. Proofs of Lemmas 9, 10, and 11
Case 1: A sandwich of the form $ded$ blocks $a$ in $X$. Note that since $d\notin R^{\prime}$, we have $de\subseteq S^{\prime}$. So, if neither $dat\subseteq S^{\prime}$ nor $adt\subseteq S^{\prime}$, then $atde\subseteq S^{\prime}$ and $e\neq t$. So to show that $X$ satisfies (i), we only need to show that $e\notin P$. First note that $e\neq a$, since $ded$ blocks $a$. If $e=r$, then, since Claim 19 implies that $art\subseteq P$, we have $X\supseteq S^{\prime}dP\supseteq atdrdart\cong abdtdatb\cong C_{2,5}$, a contradiction. We get a similar contradiction if $P=arbrt$ or $P=arcbct$ and $e=b$, since in either case $abt\subseteq P$, so $S^{\prime}dP\supseteq atdbdabt\cong abdtdatb\cong C_{2,5}$, a contradiction. If $P=arcbct$ and $e=c$, then $act\subseteq P$, so $S^{\prime}dP\supseteq atdcdact\cong abdtdatb\cong C_{2,5}$, a contradiction. With Claim 19, this completes the proof of this case.
@@
Case 2.1: $e=t$. In this case, $ta=ea\subseteq R^{\prime}$ and, by Claim 20 a sandwich of the form $trt$ does not block $a$ in $X=S^{\prime}dR^{\prime}$. The fact that $at\subseteq S^{\prime}$ then implies that $rta\nsubseteq R^{\prime}$. Therefore, there exists a $t$ in $R^{\prime}$ that is before both an $a$ in $R^{\prime}$ and the first $r$ in $R^{\prime}$. So, if $P=arat$, then we have either $tarat=eP$ or $atrat$ in $R^{\prime}$, so (ii) holds. Furthermore, if $P=rart$, we have $trart=eP$ in $R^{\prime}$, so (ii) holds. If $P=arbrt$, then one option is $tarbrt=eP\subseteq R^{\prime}$, and (ii) holds in this case. By Claim 20, there is no $rbr$ sandwich that blocks $a$, so the only other possibilities when $P=arbrt$ are $atrabrt\subseteq R^{\prime}$ or $atrbart\subseteq R^{\prime}$, and, in either case, we have $trart\subseteq R^{\prime}$, so (ii) holds. Finally, assume $P=arcbct$. If $tarcbct=eP\subseteq R^{\prime}$, then $tacbct\subseteq R^{\prime}$, so (ii) holds. Otherwise, because there is no $cbc$ sandwich that blocks $a$, at least one of $atracbct\supseteq tacbct$, $atrcabct\supseteq tcact$, or $atrcbact\supseteq tcact$ is a subordering of $R^{\prime}$, so (ii) holds.
Case 2.2: $e=r$. In this case, $r\in S^{\prime}$ and $ra=ea\subseteq R^{\prime}$. Recall that, by Claim 20, there is no sandwich of the form $rtr$ that blocks $a$ in $X$. Therefore, $r$ must follow $at$ in $S^{\prime}$, so $atr\subseteq S^{\prime}$. Because $art\subseteq P$ and $a$ is blocked by a $d$-sandwich, Lemma 10 (with $r$ playing the role of $b$) implies that a subpattern of $X$ is in $\Gamma$, a contradiction.
@@
Assume first that $e\in R$ and recall that $r$ was defined to be the first color in $R$ that is neither $t$ nor $a$. Therefore, $rea\subseteq R^{\prime}$, which implies that a sandwich of the form $ere$ blocks $a$ in $X\supseteq S^{\prime}dR^{\prime}$. Note that this, with Claim 20, implies that $e\notin P$. Recall that Claim 20 implies that every color in $atP$ is unsortable except $a$. If $e$ is blocked by a sandwich consisting of colors only in $P$, then the subordering formed by removing all colors except those in $P$ and $e$ from $X-d$ is trivially unsortable. This contradicts the assumption that all proper suborderings of $X$ are sortable.
Now assume that $e\notin R$. In this case, $P\subseteq R$ implies that $e\notin P$. By the definition of $R^{\prime}$, the fact that $e\in R^{\prime}$ implies that the last $e$ follows the last $d$ in $X$. Since $e\notin R$ and the state after $d$ is sorted is $(S,R)$, the color $e$ was sorted before the color $d$. Since $d$ was sorted before any color in $P$, the color $e$ is not blocked by a sandwich consisting only of colors in $P$ in $X$.
So, in both cases, $e\notin P$ and $e$ is not blocked by a sandwich consisting only of colors in $P$ in $X$. In particular, the subordering $ata$ does not block $e$ in $X$. This means that $e$ does not follow $a$ in $R^{\prime}$. Note first that if $eP\subseteq R^{\prime}$, then (iii) holds. So we can assume $eP\nsubseteq R^{\prime}$. This, with the fact that $e$ does does not follow $a$ in $R^{\prime}$ and $e\in R^{\prime}$, implies that $P=rart$ and $reart\subseteq R^{\prime}$. Therefore, at least one of $eatdreart$, $aetdreart$, or $atedreart$ is a subordering of $X=S^{\prime}dR^{\prime}$. This is a contradiction, because $eatdreart\supseteq eatrear\cong dabtdat\cong I_{1,0}$, $aetdreart\supseteq aetreart\cong adbtdatb\cong\mathcal{C}_{2,4}$, and $atedreart\supseteq atereart\cong abdtdatb\cong\mathcal{C}_{2,5}$. ∎
When Claim 21(i) holds, one of the orderings in the following table is a subordering of $X$.
@@
If Claim 21(iii) holds, then by swapping $d$ and $e$, we can see that one of orderings in the preceding table is a subpattern of $X$. Finally, if Claim 21(ii) holds, then one of the orderings in the following table is a subpattern of $X$.
@@
3. Further Directions
While the process of foot sorting lends itself naturally to sorting with a stack, one could consider analogs of the sock sorting problem which utilize other stack configurations or other data structures entirely.
3.1. Sorting with Multiple Stacks
Defant and Kravitz introduced a notion of foot-sorting with multiple feet in [1], though with more stacks in series, the sorting process becomes increasingly computationally intensive. It would be interesting to investigate perhaps the simplest extension of our work, the basis of sock orderings which are foot-sortable with two stacks in series.
Defant and Kravitz proved in [1] that for any number $n\geq 1$, one may construct a sock ordering which is unsortable with $n$ stacks. Using this proof, one can check that the sock ordering of the form $(abcde)^{1440}$ (that is, the block $abcde$, where $a,b,c,d$ and $e$ are distinct sock colors, repeated 1440 times) is unsortable with two stacks in series, but it might be challenging to construct a single sock ordering that is minimally unsortable in this setting.
3.2. Deque Sorting
Because a deque, or double-ended queue, allows for more access to the socks for addition and removal, one might think about deque-sorting for socks independently of a foot; a sock could be placed either over an outermost sock or within an innermost sock, and the same is possible for sock removal.111 Using a deque in the context of permutation sorting has been consider previously (e.g. see [3], [2], and [4]). This modifies the original operations of foot sorting, using the language of the deque, to be
  1. ($\text{stack}_{\uparrow}$)
    Take the leftmost sock on the right and place it on top of the deque (potentially over any socks that are already in the deque).
  2. ($\text{stack}_{\downarrow}$)
    Take the leftmost sock on the right and place it on the bottom of the deque (potentially under any socks that are already in the deque).
  3. ($\text{unstack}_{\uparrow}$)
    Remove the topmost sock from the deque and place it so that it becomes the rightmost sock on the left.
  4. ($\text{unstack}_{\downarrow}$)
    Remove the bottommost sock from the deque and place it so that it becomes the rightmost sock on the left.
While the addition of possible moves in this scenario makes for more difficult computation, some similar ideas hold as in the stack setting: in particular, it is now necessary to avoid the pattern $abab$ from appearing in the deque, just as it is necessary to avoid $aba$ appearing in the stack. It would be interesting to enumerate the basis of deque-sortable sock orderings, just as we have enumerated the basis of foot-sortable sock orderings.
It is straightforward to show that the basis of deque-sortable sock orderings is different from $\Gamma$, as well as to show that this basis is infinite. Very closely mirroring the proof of Observation 2 under the relaxed restrictions of the deque, one can generate the class of sock orderings of the form $dabtI_{1,n}$ for any $n\geq 0$ and verify that such sock orderings are minimally unsortable with the deque. However, with the increased computational difficulty in the deque setting, we do not know whether other minimal deque-unsortable sock orderings may be generated from $\Gamma$ in a similar fashion.